OpenStudy (bbb911):

This is Geometry: square root of 4/11

OpenStudy (anonymous):

0.6031

OpenStudy (angela210793):

\[\sqrt{4/11}=2/\sqrt{11}=2*\sqrt{11}/11\]

OpenStudy (anonymous):

.6030

OpenStudy (amistre64):

sqrt(4/11) is geometry?

OpenStudy (anonymous):

0.6030 or 0.603

OpenStudy (bbb911):

and in radial form?

OpenStudy (bbb911):

yes it is

OpenStudy (bbb911):

does anyone know the radial form of 0.603?

OpenStudy (amistre64):

cant say that I ever even heard of a radial form

OpenStudy (bbb911):

oh i mean radical form

OpenStudy (amistre64):

\[\sqrt{{4\over11}}=\frac{\sqrt{4}}{\sqrt{11}}=\frac{2}{\sqrt{11}}\]

OpenStudy (amistre64):

\[\frac{2}{\sqrt{11}}*\frac{\sqrt{11}}{\sqrt{11}}=\frac{2\sqrt{11}}{11}\]

OpenStudy (bbb911):

Thank you so much :D

OpenStudy (amistre64):

youre welcome :)

OpenStudy (angela210793):

Haven't I written the same????????

OpenStudy (bbb911):

oh sorry i didn't see it because of so many poeple anwsering it but i have gave you a medal to thank you :D

OpenStudy (amistre64):

angela dint use proper use of paranthesis ;)

OpenStudy (angela210793):

Oops...sorry didn't mean to ''scare'' you :):) :P